OpenStudy (anonymous):

Evaluate y = x – 13 when x = –3.

pooja195 (pooja195):

Plug in the number y=-3-13 -3-13=?

OpenStudy (anonymous):

=y?

pooja195 (pooja195):

no -3-13 is what? what does it equal?

OpenStudy (anonymous):

-16

pooja195 (pooja195):

Good so y=?

OpenStudy (anonymous):

-16

pooja195 (pooja195):

good ^.^

OpenStudy (anonymous):

Thank you! I got one more could you help me with

OpenStudy (anonymous):

Solve the system of equations. 2x+3y=1 y=3x+15 ( , )

pooja195 (pooja195):

Sub in y in equation 1 2x+3(3x+15)=1 solve for x

OpenStudy (anonymous):

?

pooja195 (pooja195):

2x+3y=1 <--equation 1 y=3x+15 <---equation 2 Using substituion we take equation 2 and plug it into equation 1 because we know the value of y so 2x+3(3x+15)=1 from here solve for y